# reference.bib

### v1

@@ 8. Theta Correspondence
$\displaystyle=\int_{\textup{Sym}_{2}(\mathbb{R}/\mathbb{Z})}\int_{\mathcal{R}(2,\mathbb{R})}(\det Y)^{(s+1)/2-r}\overline{R[\delta_{k}^{(r)}\Theta](Z,W)}\hbox{d}^{*}Z,$
where in the third equation, we used the usual unfolding trick. Here, $C(2,\mathbb{R})+i\mathcal{R}(2,\mathbb{R})$ is a fundamental domain for the action of $P_{2,0}$ on $\mathbb{H}_{2}$, where $C(2,\mathbb{R})$ denotes a fundamental parralepiped of $\textup{Sym}(2,\mathbb{Z})$ in $\textup{Sym}(2,\mathbb{R})$ and $\mathcal{R}(2,\mathbb{R})$ the Minkowski reduced matrices, as in [Krieg_Book, Proposition 2.2, page 146].
Now, from Proposition 7.1 and Lemma 7.1, we have
$\int_{\textup{Sym}_{2}(\mathbb{R}/\mathbb{Z})}R[\delta_{k}^{(r)}\Theta](Z,W)\hbox{d}X=\sum_{\begin{subarray}{c}\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})\\ \textup{rank}\ell=2,S_{1}[\ell]=0\end{subarray}}R_{0}^{*}\left[p\left(\det(R_{W}[\ell]Y),\textup{tr}\left(R_{W}[\ell]Y\right)\right)e^{-\pi\textup{tr}\left(R_{W}[\ell]Y\right)}\right],$
where $R_{0}^{*}=(\det Y)^{2-r}\det(\partial_{Y})(\det Y)^{1+2r}\det(\partial_{Y})$. The fact that $S_{1}[\ell]=0$ follows from the first statement of Lemma 7.1. The rest of the expression of the integral follows from the second part of that lemma and from the way $R$ has been defined. Hence, from [Krieg_Book, Proposition 4.4, 4.5], the integral becomes
$I=\sum_{\begin{subarray}{c}\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})/\textup{GL}_{2}(\mathbb{Z})\\ \textup{rank}\ell=2,S_{1}[\ell]=0\end{subarray}}\int_{\mathcal{P}_{2}}(\det Y)^{s/2}R_{0}\left[p\left(\det(R_{W}[\ell]Y),\textup{tr}\left(R_{W}[\ell]Y\right)\right)e^{-\pi\textup{tr}\left(R_{W}[\ell]Y\right)}\right]\hbox{d}^{*}Y,$
where now $R_{0}=(\det Y)^{-(1+r)}R_{0}^{*}$ and $\hbox{d}^{*}Y=(\det Y)^{-\frac{3}{2}}\hbox{d}Y$ is the invariant measure for the action of $\textup{GL}_{2}$ on $\mathcal{P}_{2}$. Now, for every $\ell$ in our sum, we have that $R_{W}[\ell]$, hence $(R_{W}[\ell])^{-1}$, is positive definite. So, we can write $(R_{W}[\ell])^{-1}=AA^{t}$ with $A$ lower triangular and change variables $Y\longmapsto Y[A^{t}]$. Then $\det(R_{W}[\ell]Y)$ and $\textup{tr}\left(R_{W}[\ell]Y\right)$ become $\det Y$ and tr $\textup{tr }Y$ respectively and the measure remains invariant. Hence, since $R_{0}$ is $\textup{GL}_{2}$-invariant (cf. [bocherer, (2.4)]), the integral becomes
@@
where $\phi_{2}(t)=t\left(t-\frac{1}{2}\right)$ (cf. [maass, page 82]). Therefore, we have:
$\displaystyle I$ $\displaystyle=\sum_{\begin{subarray}{c}\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})/\textup{GL}_{2}(\mathbb{Z})\\ \textup{rank }\ell=2,S_{1}[\ell]=0\end{subarray}}(\det R_{W}[\ell])^{-s/2}\int_{\mathcal{P}_{2}}(\det Y)^{s/2-2r}M\left[(\det Y)^{1+2r}\det(\partial_{Y})\left[p(\det Y,\textup{tr}Y)e^{-\pi\textup{tr}Y}\right]\right]\hbox{d}^{*}Y$
@@
$\displaystyle=\sum_{\begin{subarray}{c}\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})/\textup{GL}_{2}(\mathbb{Z})\\ \textup{rank }\ell=2,S_{1}[\ell]=0\end{subarray}}(\det R_{W}[\ell])^{-s/2}\phi_{2}(s/2-2r)\phi_{2}(s)\int_{\mathcal{P}_{2}}(\det Y)^{s/2}p(\det Y,\textup{tr}Y)e^{-\pi\textup{tr}Y}\hbox{d}^{*}Y$
@@
where the second and third equality follow after transferring $M$ to its adjoint. The last equality follows from Proposition 5.1 and Lemma 5.5. We now just need to compute the last integral. It is true that (cf. [maass])
@@
for any $s\in\mathbb{C}$ with $\textup{Re}(s)>1/2$ and $T\in\mathcal{P}_{2}$. Here $\Gamma_{2}(s):=\Gamma(s)\Gamma\left(s-1/2\right)$. Setting now $T\longmapsto\pi T$ and then applying the operator $\delta_{1}^{(r)}(S,T)$, where $S$ is symmetric, we obtain from Lemma 7.1
$\int_{\mathcal{P}_{2}}(\det Y)^{s/2}(\det T)^{-3r-1}p(\det(YT),\textup{ tr}(YT))e^{-\pi\textup{tr}(YT)}\hbox{d}^{*}Y=\pi^{1/2-s}\Gamma_{2}(s/2)\delta_{1}^{(r)}[(\det T)^{-s/2}].$
But for any $\alpha\in\mathbb{Z}$, we have
$\displaystyle\delta_{\alpha}\left[(\det T)^{-s/2}\right]$ $\displaystyle=(\det T)^{1/2-\alpha}\det(\partial_{U})(\det T)^{\alpha-1/2-s/2}$
$\displaystyle=-\frac{1}{4}(\det T)^{1/2-\alpha}\det(\partial_{T})(\det T)^{\alpha-1/2-s/2}$
$\displaystyle=-\frac{1}{4}(\det T)^{1/2-\alpha}(\alpha-1/2-s/2)(\alpha-s/2)(\det T)^{\alpha-3/2-s/2}$
$\displaystyle=-\frac{1}{4}\phi_{2}((s+1)/2-\alpha)(\det T)^{-s/2-1},$
because $\det(\partial_{U})=-\det(\partial_{T})/4$ and the third equality follows from a well-known formula (see [cayley_identity, Theorem 2.2] for example). Hence
$\delta_{1}^{(r)}\left[(\det T)^{-s/2}\right]=(-4)^{r}\prod_{j=1}^{r}\phi_{2}(s/2-2j+3/2)(\det T)^{-s/2-r}.$
@@
Corollary 1.
From Proposition 4.2 and Theorem 8.1, we obtain the meromorphic continuation of $\mathcal{D}_{F,G}$ to $\mathbb{C}$, due to the one of $\widetilde{E}(Z,s)$, as we noted in the beginning of section 8.
Remark 2.
The conditions of 8.1 are satisfied for at least the $A_{4},D_{4}$ and $E_{8}$ lattices. These are Euclidean in the sense of [eisenstein_thesis, Section 1.6.4]). Also, for each of these lattices, their Grammian matrix is known (see [eisenstein_thesis, Section 1.6.2]) and therefore if we take $x$ to be the first standard basis vector, we have $S[x]=2$. Hence, from Remark 1, the condition #$\#\mathcal{C}^{1}(\Gamma_{S})=1$ is true in our setting.
@@ 9. The $E_{8}$ lattice
As an application, we obtain a precise result regarding the functional equation of $\mathcal{D}_{F,G}$ in the case when
@@
This is a positive definite, even matrix with $\det S=1$ and the lattice it corresponds to is the so-called $E_{8}$ lattice (cf. [eisenstein_thesis, Example 1.2.10]). This is the unique unimodular lattice with only one-dimensional cusp ([eisenstein_thesis, Example 1.6.20]). As this is unimodular, the level $q$ is 11 11 and so the character $\chi_{S}$ of Proposition 6.1 is trivial. Therefore, in this case, $\Gamma_{0}^{(2)}(q)$ is the whole $\hbox{Sp}_{2}(\mathbb{Z})$ and the symplectic Eisenstein series of section 8 is
$\widetilde{E}(Z,s)=\sum_{\gamma\in P_{2,0}\backslash\textup{Sp}_{2}(\mathbb{Z})}(\det\hbox{Im}(\gamma Z))^{s}.$
Now, if $\xi(s)=\pi^{-s/2}\Gamma(s/2)\zeta(s)$, it is well known (see for example [kalinin, Theorem 2]), that the modified Eisenstein series
~$\widetilde{\mathcal{E}}(Z,s):=\xi(2s)\xi(4s-2)\widetilde{E}(Z,s)$
has a meromorphic continuation to $\mathbb{C}$ with a simple pole at $s=3/2$ and satisfies the functional equation
~$\widetilde{\mathcal{E}}\left(Z,3/2-s\right)=\widetilde{\mathcal{E}}(Z,s).$
@@
$(4\pi)^{-s}\Gamma(s)D_{F,G}(s)=\#\hbox{SO}(S;\mathbb{Z})\cdot\langle F(Z)\cdot E(Z,s-k+9)\textup{, }G(Z)\rangle.$
@@
$D_{F,G}^{*}(s):=(4\pi)^{-s}\Gamma(s)\xi(s-k+6)\xi(2s-2k+10)\xi(s-k+9)\xi(s-k+8)\gamma_{S}(s-k+9)D_{F,G}(s),$
@@
$D_{F,G}^{*}(s)=\#\hbox{SO}(S;\mathbb{Z})\cdot\langle F(Z)\cdot E^{*}(Z,s-k+9)\textup{, }G(Z)\rangle.$
@@
Let $S$ be as above, corresponding to the $E_{8}$ lattice. With the notation as above, $D_{F,G}^{*}(s)$ has a meromorphic continuation to $\mathbb{C}$ and is invariant under $s\longmapsto 2k-9-s$.
[…]

### v2

@@ 8. Theta Correspondence
$\displaystyle=\int_{\mathcal{C}(2,\mathbb{R})}\int_{\mathcal{R}(2,\mathbb{R})}(\det Y)^{(s+1)/2-r}\overline{R[\delta_{k}^{(r)}\Theta](Z,W)}\hbox{d}^{*}Z,$
where in the second equation we used the invariance up to $\chi$ of $R[\delta_{k}^{(r)}\Theta]$ and in the third equation we used the usual unfolding trick. Here, $C(2,\mathbb{R})+i\mathcal{R}(2,\mathbb{R})$ is a fundamental domain for the action of $P_{2,0}$ on $\mathbb{H}_{2}$, where $C(2,\mathbb{R})$ denotes a fundamental parralepiped of $\textup{Sym}(2,\mathbb{Z})$ in $\textup{Sym}(2,\mathbb{R})$ and $\mathcal{R}(2,\mathbb{R})$ the Minkowski reduced matrices, as in [Krieg_Book, p. 29]. In the following, we write $\textup{Sym}_{2}(\mathbb{R}/\mathbb{Z})$ for $\mathcal{C}(2,\mathbb{R})$. Now, from Proposition 7.2, Lemma 7.1 and the proof of Lemma 7.3, we have
$\int_{\textup{Sym}_{2}(\mathbb{R}/\mathbb{Z})}\overline{R[\delta_{k}^{(r)}\Theta](Z,W)}\hbox{d}X=\\ =\sum_{\begin{subarray}{c}\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})\\ \textup{rank }\ell=2,S_{1}[\ell]=0\end{subarray}}R_{0}^{*}\left[p\left(\det(R_{W}[\ell]Y),\textup{tr}\left(R_{W}[\ell]Y\right)\right)e^{-\pi\textup{tr}\left(R_{W}[\ell]Y\right)}\right],$
where $R_{0}^{*}=(\det Y)^{2-r}\det(\partial_{Y})(\det Y)^{1+2r}\det(\partial_{Y})$. The fact that only terms with $S_{1}[\ell]=0$ remain follows from (19). The rest of the expression of this integral follows from (14) and from the fact that $R=(\phi^{*})^{-1}\left(R_{0}^{*}(\det Y)^{-(1+r)}\right)$ (note that from Lemma 7.1, the polynomial $p$ has real coefficients). Hence, from [Krieg_Book, Proposition I.4.4, I.4.5], the integral $I$ becomes
$\sum_{\begin{subarray}{c}\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})/\textup{GL}_{2}(\mathbb{Z})\\ \textup{rank}\ell=2,S_{1}[\ell]=0\end{subarray}}\int_{\mathcal{P}_{2}}(\det Y)^{s/2}R_{0}\left[p\left(\det(R_{W}[\ell]Y),\textup{tr}\left(R_{W}[\ell]Y\right)\right)e^{-\pi\textup{tr}\left(R_{W}[\ell]Y\right)}\right]\hbox{d}^{*}Y,$
where now $R_{0}=(\det Y)^{-(1+r)}R_{0}^{*}$ and $\hbox{d}^{*}Y=(\det Y)^{-\frac{3}{2}}\hbox{d}Y$ is the invariant measure for the action of $\textup{GL}_{2}$ on $\mathcal{P}_{2}$. Now, for every $\ell$ in our sum, we have that $R_{W}[\ell]$, hence $(R_{W}[\ell])^{-1}$, is positive definite. So, we can write $(R_{W}[\ell])^{-1}=AA^{t}$ with $A$ lower triangular and change variables $Y\longmapsto Y[A^{t}]$. Then $\det(R_{W}[\ell]Y)$ and $\textup{tr}\left(R_{W}[\ell]Y\right)$ become $\det Y$ and $\textup{tr }Y$ respectively and the measure remains invariant. Now, from [bocherer, equation (2.4)], $R_{0}$ is $\textup{GL}_{2}$-invariant ($n=2$ and $m=2+2r$ in the notation there). Hence, the integral becomes
@@
where $\displaystyle{\phi_{2}(t)=t\left(t-\frac{1}{2}\right)}$ (see [bocherer, (3.3)]). Therefore, we have
[…]
@@
$\displaystyle=\sum_{\begin{subarray}{c}\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})/\textup{GL}_{2}(\mathbb{Z})\\ \textup{rank }\ell=2,\ S_{1}[\ell]=0\end{subarray}}(\det R_{W}[\ell])^{-s/2}\phi_{2}(s/2-2r)\phi_{2}(s/2)\int_{\mathcal{P}_{2}}(\det Y)^{s/2}p(\det Y,\textup{tr}Y)e^{-\pi\textup{tr}Y}\hbox{d}^{*}Y$
@@
where the second and third equality follow after transferring $M$ to its adjoint. The last equality follows from Proposition 5.3 and Lemma 5.10. We now just need to compute the last integral. It is true that (cf. [maass, page 80, 81])
@@
for any $s\in\mathbb{C}$ with $\textup{Re}(s)>1/2$ and $T\in\mathcal{P}_{2}$. Here $\Gamma_{2}(s):=\Gamma(s)\Gamma\left(s-1/2\right)$. Setting now $T\longmapsto\pi T$ and then applying the operator $\delta_{1}^{(r)}(S,T)$, where $S$ is symmetric, so that $U:=S+iT\in\mathbb{H}_{2}$, we obtain from Lemma 7.1 and (14)
$\int_{\mathcal{P}_{2}}(\det Y)^{s/2}(\det T)^{-r}p(\det(YT),\textup{ tr}(YT))e^{-\pi\textup{tr}(YT)}\hbox{d}^{*}Y=\pi^{1/2-s}\Gamma_{2}(s/2)\delta_{1}^{(r)}[(\det T)^{-s/2}].$
But for any $\alpha\in\mathbb{Z}$ and $w\in\mathbb{C}$, we have from (11)
$\displaystyle\delta_{\alpha}\left[(\det T)^{w}\right]$ $\displaystyle=(\det T)^{1/2-\alpha}\det(\partial_{U})(\det T)^{\alpha-1/2+w}$
$\displaystyle=-\frac{1}{4}(\det T)^{1/2-\alpha}\det(\partial_{T})(\det T)^{\alpha-1/2+w}$
$\displaystyle=-\frac{1}{4}(\det T)^{1/2-\alpha}(\alpha-1/2+w)(\alpha+w)(\det T)^{\alpha-3/2+w}$
$\displaystyle=-\frac{1}{4}\phi_{2}(w+\alpha)(\det T)^{w-1},$
because $\det(\partial_{U})B(T)=-\det(\partial_{T})B(T)/4$ for a function $B=B(T)$ depending only on $T$ and the third equality follows from a well-known formula (see [cayley_identity, Theorem 2.2] for example). Hence, by successively applying the above and using the fact that $\phi_{2}(s)=\phi_{2}(1/2-s)$ for any $s\in\mathbb{C}$, we get
$\delta_{1}^{(r)}\left[(\det T)^{-s/2}\right]=(-4)^{-r}\prod_{j=1}^{r}\phi_{2}\left(\frac{s-(2j-1)}{2}\right)(\det T)^{-s/2-r}.$
@@
Corollary 8.3.
Assume $S$ satisfies the assumptions of Theorem 8.2. Then, $E(W,s)$ admits a meromorphic continuation to the complex plane and
$\Gamma\left(\frac{s+1}{2}-r\right)\Gamma(s-2r)L_{q}(s+1-2r,\chi)\zeta_{q}(2s-4r)\xi(s)\xi(s-1)\gamma_{S}(s)E(W,s)$
has only possible simple poles at $s\in\{(n+2)/2,\ (n+4)/2\}$ if $q\neq 1$ and at $s\in\{(n-2)/2,\ n/2,\ (n+2)/2,\ (n+4)/2\}$ if $q=1$. Here, $L(s,\chi)$ denotes the usual Dirichlet $L$-function attached to $\chi$, and the subscript q means that we omit the Euler factors sharing prime factors with $q$.
Proof.
From [Shimura_Euler_Product, Theorem 19.3], we have that
$\Gamma(s)\Gamma(2s-1)L_{q}(2s,\chi)\zeta_{q}(4s-2)\widetilde{E}(Z,\chi,s)$
has a meromorphic continuation to the complex plane with possible simple poles only at $s\in\{1,3/2\}$ if $q\neq 1$ and at $s\in\{0,1/2,1,3/2\}$ if $q=1$ (see [Shimura_Euler_Product, (19.3.1), (19.3.2)] because the character $\chi$ satisfies $\chi^{2}=1$). Hence, the Corollary follows from Theorem 8.2. ∎
Corollary 8.4.
From Proposition 4.4 and Corollary 8.3, we obtain the meromorphic continuation of $\mathcal{D}_{F,G}(s)$ to $\mathbb{C}$, due to the one of $\widetilde{E}(Z,\chi,s)$, as we noted in the beginning of Section 8.
Remark 8.5.
The conditions of Theorem 8.2 are satisfied when $S$ corresponds to the matrices defining the quadratic forms in $4$ and $8$ variables listed in the tables of [hanke2011enumerating, Section 8]. These have one class in the genus of maximal lattices, and by looking at their Grammian matrix, if $x$ is the first standard basis vector, we have $S[x]=2$. Hence, from Remark 5.9, the condition $\#\mathcal{C}^{1}(\Gamma_{S})=1$ is true in this setting.
@@ 9. The $E_{8}$ lattice
As an application, we obtain a precise result regarding the functional equation of $\mathcal{D}_{F,G}(s)$ in the case when
@@
This is a positive definite, even matrix with $\det S=1$ and the lattice it corresponds to is the so-called $E_{8}$ lattice (cf. [eisenstein_thesis, Example 1.2.10]). This is the unique unimodular lattice with only one-dimensional cusp ([eisenstein_thesis, Example 1.6.20]). As this is unimodular, the level $q$ is $1$ and so the character $\chi_{S}$ of Proposition 6.5 is trivial. Therefore, in this case, $\Gamma_{0}^{(2)}(q)$ is the whole $\textup{Sp}_{2}(\mathbb{Z})$ and the symplectic Eisenstein series of (20) is
$\widetilde{E}(Z,s)=\sum_{\gamma\in P_{2,0}\backslash\textup{Sp}_{2}(\mathbb{Z})}(\det\textup{Im}(\gamma Z))^{s}.$
Now, if $\xi(s)=\pi^{-s/2}\Gamma(s/2)\zeta(s)$, from [Shimura_Euler_Product, Theorem 19.3] (as in Corollary 8.3), the modified Eisenstein series
$\widetilde{\mathcal{E}}(Z,s):=\xi(2s)\xi(4s-2)\widetilde{E}(Z,s)$
has a meromorphic continuation to $\mathbb{C}$ with possible simple poles only at $s\in\{0,1/2,1,3/2\}$, and from [kalinin, Theorem 2] satisfies the functional equation
$\widetilde{\mathcal{E}}\left(Z,3/2-s\right)=\widetilde{\mathcal{E}}(Z,s).$
@@
$(4\pi)^{-s}\Gamma(s)\mathcal{D}_{F,G}(s)=\#\textup{SO}(S;\mathbb{Z})\cdot\langle F(W)\cdot E(W,s-k+9)\textup{, }G(W)\rangle.$
@@
$\mathcal{D}_{F,G}^{*}(s):=(4\pi)^{-s}\Gamma(s)\xi(s-k+6)\xi(2s-2k+10)\xi(s-k+9)\xi(s-k+8)\gamma_{S}(s-k+9)\mathcal{D}_{F,G}(s),$
@@
$\mathcal{D}_{F,G}^{*}(s)=\#\textup{SO}(S;\mathbb{Z})\cdot\langle F(W)\cdot E^{*}(W,s-k+9)\textup{, }G(W)\rangle.$
@@
Let $S$ be as above, corresponding to the $E_{8}$ lattice. With the notation as above, $\mathcal{D}_{F,G}^{*}(s)$ has a meromorphic continuation to $\mathbb{C}$ and is invariant under $s\longmapsto 2k-9-s$.
Data Availability: Data sharing not applicable to this article as no datasets were generated or analysed during the current study.